OpenStudy (anonymous):

for a circle of radius 9 feet, find the arc length subtended by a central angle of 52 degrees? A. 13pi/5 b. 11pi/5 c. 13pi/11 d. 5pi/11

OpenStudy (anonymous):

Arclength = \[\Theta*r\]

OpenStudy (anonymous):

how do you do that

OpenStudy (campbell_st):

well 1st you need to change 52 degrees to radians... to use the formula... can you don that....?

OpenStudy (anonymous):

0.91?

OpenStudy (campbell_st):

well leave it in terms of pi \[52^o = \frac{52 \pi}{180}\] then use the formula shown above \[l = r \times \theta\] i = arc length, r = radius and theta= angle subtended... hope it helps simplify the fraction answer to get an answer matches one of the choices

OpenStudy (anonymous):

how do you time it by theta though?

OpenStudy (campbell_st):

well look at it this way \[l = 9 \times \frac{52\pi}{180}\] cancel the common factor of 9 so \[l = \frac{52\pi}{20} \] just finsih cancelling common factors.

OpenStudy (anonymous):

ok i understand now thankyou